OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9 over 5 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is.

OpenStudy (whpalmer4):

\[C = \frac{9}{5}(F-32)\]If you want to solve that for F, I would multiply both sides by 5. Then expand the right hand side by the distributive property. Add or subtract any numbers from both sides to give you an F term alone on one side. If F has a number in front of it (a coefficient), divide both sides of the equation by that number. As a test, you can use -40, which is the point on both scales where the temperature is the same! Part 2 is just using your formula with C = 6.

OpenStudy (anonymous):

\[C=\frac{ 9 }{5 }\left( F-32 \right)\] multiply by 5/9 \[\frac{ 5 }{ 9 }C=\frac{ 5 }{9 }*\frac{ 9 }{5 }\left( F-32 \right)\] \[\frac{ 5 }{ 9 }C=F-32\] Now add 32 to both sides and get the required solution. for part 2,put c=6 in the given formula.